OpenStudy (anonymous):

Simplify the radical expression sqrt 150x^3k^4 a)75x^3k^2 b)5xk times sqrt 6xk c)5xk^2 times sqrt 6x d)5 times sqrt 6x^3k^4 Please help! I tried asking people I know how to solve these types of problem but they tell me to look in the textbook. The textbook explains these problems so confusing! I'll be so so so grateful if someone out there can help me with this problem and a few others.

Nnesha (nnesha):

\[\huge\rm \sqrt{150x^3k^4}\] is it your question ?

Nnesha (nnesha):

factor 150

Nnesha (nnesha):

x^3 can be written as \[\huge\rm \sqrt{x^3} =\sqrt{x \times x \times x } = \sqrt{x^2 \times x}\] same with k^4

OpenStudy (anonymous):

yeah \[\sqrt{150}x^3k^4\] is my question. The number's factors are 30 and 5, but I'm not sure how to factor squares. okay.

Nnesha (nnesha):

alright there are more factors of 150 one of number should be perfect square

OpenStudy (anonymous):

*square roots

Nnesha (nnesha):

yeah square roots**

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha alright there are more factors of 150 one of number should be perfect square \(\color{blue}{\text{End of Quote}}\) perfect square roots

OpenStudy (anonymous):

25 is a perfect square factor of 150 so should I use that?

Nnesha (nnesha):

yes 25 times 6 =150 so \[\sqrt{25 \times 6} = 5 \sqrt{6}\] square root of 25 is 5

Nnesha (nnesha):

x^3 can be written as \[\huge\rm \sqrt{x^3} =\sqrt{x \times x \times x } = \sqrt{x^2 \times x}\] same with k^4 did you get this one ?

Nnesha (nnesha):

remember when you multiply same bases you should add their exponent so x^3 can be written as x^2 times x i wrote x to the 2nd power so you can cancel square root with square

OpenStudy (anonymous):

Yeah I remember that. So I factor the x and k too? I see what you are saying but how do I write that with the factored number?

Nnesha (nnesha):

yes right k^4 can be written as k^2 times k^2 then you can cancle square root with squares

Nnesha (nnesha):

 like this